what are the solutions of the quadratic equation 3xsquared+5x=-4
Answer by dabanfield(803)

Are you sure the quadratic is
3x^2 + 5x = -4 or
3x^2 + 5x + 4 = 0
There are no REAL solutions to this equation. They are both imaginary (that is, of the form a+bi where i = sqrt(-1)).